OpenStudy (anonymous):

vertex of f(x)=(x+3)^2

terenzreignz (terenzreignz):

Well, you're in luck When you have your parabola in this form f(x) = (x - h)² + k Then your vertex is at the point (h , k)

OpenStudy (anonymous):

Yea but what are the numbers?

terenzreignz (terenzreignz):

What indeed... ;) Observe it more carefully... f(x) = (x+3)² f(x) = (x-h)² + k So what are h and k?

OpenStudy (anonymous):

-3, 0?

terenzreignz (terenzreignz):

There you go ^.^